# A Survey on Human-AI Teaming with Large Pre-Trained Models

### v1

@@ 2.1.1 Human-AI Collaboration with Active Learning
Unlike traditional machine learning paradigms, active learning involves an iterative process where a model selectively identifies data requiring labeling, optimizing system performance with minimal training data [47]. Particularly, this approach is crucial for fine-tuning large pre-trained models, especially transformer-based ones, which require substantial labeled data for specific user scenarios. Active learning efficiently harnesses human expertise to pinpoint areas of uncertainty, enabling more targeted training with less annotation effort [138]. The active learning workflow begins with an unlabeled dataset and a pre-trained model. The model predicts labels for each sample, outputting confidence levels. When predictions fall below a quality threshold, human annotators step in for manual annotation. This iterative process of re-training the model with new labeled data continues until satisfactory confidence levels are achieved. Human AI teaming plays a vital role in efficient data labeling and continuous model improvement.
@@
Recent research in this area has introduced novel methods to leverage human expertise effectively. For instance, Pries et al. [127] propose a method for efficient data labeling using precise distance measurements to filter data samples, streamlining the presentation of data to experts. Lu et al. [95] discuss human-guided interventions for continuous model improvement, focusing on dissociating biases. Bao et al. [8] explore the transformation of human-annotated rationales into continuous attention mechanisms, enhancing the learning of domain-invariant representations. Additionally, Lertvittayakumjorn et al. [90] present a generalizable approach for debugging deep text classification models with human input, applicable to larger models. Finally, Yao et al. [175] propose an explainable-generation Active Learning framework for simpler tasks, potentially improving the sampling process in data augmentation.
2.1.2 Reinforcement Learning from Human Feedback
@@
Reinforcement Learning from Human Feedback (RLHF) is a paradigm that integrates human insights into the training process [29, 67]. This approach enhances the collaboration between humans and AI systems, especially during the data control phase — a critical stage for the success of large pre-trained models. In RLHF, the human role extends beyond mere data provision to actively shaping the AI’s learning trajectory. This human-AI synergy is particularly effective in mitigating biases in training data, leading to fairer AI outcomes. Additionally, the human touch in RLHF makes AI decisions more transparent and comprehensible, thereby fostering greater trust among users. The adaptability of RLHF models to user preferences underscores the significance of human understanding in complex, contextual situations, enabling AI systems to be more aligned with ethical standards and personalization needs.
@@
However, the implementation of RLHF is not without its challenges. Ensuring the consistency and accuracy of human feedback, effectively scaling this approach for large datasets, and harmonizing human insights with algorithmic processes are areas that require careful consideration [87]. Overcoming these challenges is essential to maximize the benefits of RLHF, paving the way for AI models that are not only technically proficient but also deeply attuned to human values and needs.
@@
Despite various theories and empirical models behind RL, as mentioned in [73], RL frameworks can be abstracted as a continuous interaction between agents and their surrounding environment, in which they take actions individually or cooperatively given their states and state transitions to maximize long-term rewards. Traditionally, given a designated reward model, if the accumulated long-term reward is exactly or approximately converged under limited horizons, the agent’s behaviors and the corresponding dynamics [144], like policy $\pi(a|s)$, Q function $Q(s,a)$, etc. are more likely to converge. Only after OpenAI ’s introduction of instructGPT [121] to illustrate how a reward model trained by high-quality data from labelers can optimize the LM’s ranking of generative content did research communities realize RL’s critical role in improving LM performance. Since labelers are asked to label this data in multiple turns to yield the most suitable outputs, this interaction pattern under RL formalization is considered RLHF. Furthermore, this working pattern can be widely applied in similar scenarios within feedback-acquiring processes. In the HAI teaming model, once human instructions are quantitatively converted to a reward model by ranking or scoring, the interaction between human and AI system agents can be formalized and optimized by RLHF.
@@
A notable early work in RLHF is by Ziegler et al. [181], where language models were trained to improve text generation and summarization based on human preferences. However, this approach, limited by its reliance on a single-run process and restricted data, faced challenges in generalization. In contrast, OpenAI’s instructGPT [121] introduced a three-phase training paradigm: initial training with supervised policy via demonstration data, training a reward model with labeler feedback, and optimizing the ranking policy using RL against the reward model. This methodology provides a robust baseline and ensures that GPT models evolve with continuous human feedback, leading to the creation of more sophisticated and responsive models, such as the web browser-assisted question-answering agent by Nakano et al. [113].
Despite these advancements, RLHF remains a complex and often unstable process. To address this, Stanford’s recent research has introduced the Direct Preference Optimization (DPO) training model [130]. DPO simplifies policy training by directly mapping reward functions to optimal policies, circumventing the need for intricate reward maximization constraints. This approach not only surpasses traditional RLHF in controlling sentiment in generations but also enhances response quality in tasks like summarization and single-turn dialogue. Its simplicity in implementation and training offers a promising direction for further simplifying RLHF processes.
@@
The integration of Reinforcement Learning from Human Feedback (RLHF) into Human-AI (HAI) processes presents unique challenges due to the intricate steps and components involved, as evidenced by the work of Zhang et al. [179] and Johnson et al. [72]. Recognizing these complexities, recent research has shifted focus towards enhancing collaboration within the closely-coupled phases of HAI, with the aspiration of exploring further extensions in the future. A notable gap in current RLHF models is the communication barrier between AI and humans, highlighting the need for more advanced frameworks to bridge this divide.
A significant contribution to human-agent coordination is the instructRL model by Hu and Sadigh [65], which employs natural language instructions to align RL agents with human-preferred policies. This model leverages the embedded knowledge of pre-trained models, enabling AI agents to collaborate more effectively with humans across various tasks, including gaming, robotics, and control decision-making. InstructRL’s adaptability to different domains through knowledge transfer and meta-learning positions it as a versatile model for diverse HAI scenarios.
Research by Mirchandani et al. [109] suggests that Large Language Models (LLMs), even without additional training, can function as proficient general sequence modelers due to their in-context learning abilities. This research explores the use of LLMs in robotics, particularly for extrapolating state sequences and prompting reward-conditioned trajectories, despite challenges like latency and compute costs. Chain-of-thought prompting further enhances LLM performance in complex reasoning tasks, as shown by Wei et al. [163]. The Distillation and Retrieval of Online Corrections (DROC) system by Zha et al. [177] exemplifies the capability of LLMs to assimilate language feedback and distill knowledge from corrections, demonstrating effectiveness in novel settings.
However, McNeese et al.’s research [103] provides a counterpoint to the notion of RLHF as a universal solution. Despite the success of NeoCITIES [103] RL agents in simulated environments, the disparity between these environments and real-world complexities remains a significant hurdle. Furthermore, RL agents lack proficiency in non-verbal communication and still struggle with detailed verbal interactions, indicating a need for more sophisticated models that can fully comprehend both verbal and non-verbal human communication. These observations underscore the necessity for developing new HAI frameworks that go beyond existing RLHF models, calling for enhanced human integration to address the multifaceted nature of real-world environments.
2.2 Human Evaluation in AI
@@
2.2.1 Evaluation of Trust and Transparency
@@
Establishing trust and transparency is crucial in Human-Eval-AI systems. Trust, inherently linked with transparency and understandability, aligns closely with how users perceive an AI system’s abilities and limitations. As highlighted by Bansal et al. [6], it is essential for users to recognize the AI system’s potential for errors. This understanding enables users to develop a realistic mental model of the AI, crucial for anticipating possible failures or underperformance.
@@
Zhou et al.’s work [180] further underscores this by recommending the integration of uncertainty measures and machine learning model performance into AI-informed human decision-making processes. They advocate for a trial-to-trial approach in decision-making, where users adjust their judgment boundaries based on their experience of uncertainty and AI interactions. Their proposed Uncertainty-Performance Interface (UPI) allows users to concurrently evaluate the quality and performance of machine learning models, fostering deeper trust in AI systems. Similarly, Av et al. [5] emphasize the importance of incorporating intrinsic uncertainty into the collaborative process. They suggest a mechanism for human experts to offer feedback, either as corrections to current AI recommendations or insights on preferable aspects within the search space. This feedback is integrated into a Bayesian optimization model, aiming to align the model’s inferences with the expert’s domain-specific knowledge.
2.2.2 Approaches to Improve Human Evaluation
@@
In Human-Eval-AI, human input remains essential, especially in instructing AI agents and refining AI decision-making processes. The study by Chhibber et al. [25] showcases the benefits of integrating human teachings into AI, not only enhancing AI performance but also building trust and influencing job delegation practices. In a similar vein, Hu and Sadigh’s instructRL [65] introduces a novel form of human-AI interaction through high-level natural language instructions, thereby improving AI interpretability and alignment with human intentions. Aaquib et al.’s research [146] on mental modeling in human-robot teaming reveals that mutual understanding between humans and AI significantly boosts team performance, advocating for AI systems capable of adapting to human mental models. Complementing this, Correia et al. [34] emphasize context in AI system design, proposing personalized user interactions for more relevant human-machine team actions and information.
@@
Further exploring human evaluation, Clark et al. [30] examine how human evaluators can differentiate between texts produced by humans and models like GPT-2 and GPT-3. They suggest methods to enhance evaluation accuracy, including detailed instructions, annotated examples, and text comparisons. Ouyang et al. [121] present a different approach, fine-tuning GPT-3 into ‘InstructGPT’ using human feedback for improved user instruction alignment. Additionally, Lu et al.’s Error Analysis Prompting (EAPrompt) [96] focuses on evaluating machine translations by LLMs, leveraging Chain-of-Thought reasoning to refine LLM evaluative capabilities. Chen et al’s work shows domain specific evaluation of LLM in problem-solving [24]. Chiang and Lee [26] investigate LLMs for human-like text quality assessments in specific NLP tasks, highlighting similarities with human evaluations. Finally, Uchendu et al. [152] highlight the effectiveness of collaborative efforts in enhancing detection capabilities, offering a unique perspective on human collaboration in evaluations.
3 Effective Human-AI Joint Systems
@@
This section concentrates on achieving both efficiency and safety of the joint Human-AI (HAI) systems. The focus is on merging the strengths of different HAI systems to ensure optimal performance while maintaining the highest standards of safety.
3.1 Improving User Interfaces for Effective Teaming
@@
3.1.1 Adaptation and Personalization
@@
The integration of AI agents into human teams has historically been limited by their narrow adaptability and personalization capabilities. Traditional AI systems struggled to understand human subtleties and adapt to team dynamics, being confined to specific tasks without the ability to learn or adjust, thereby reducing their effectiveness in collaborative settings [179].
@@
The introduction of large pre-trained models has significantly revolutionized AI’s adaptability and personalization in teamwork. These models, trained on extensive datasets and utilizing sophisticated algorithms, have a markedly improved capacity to comprehend and adapt to team changes. Research demonstrates that LLM-equipped AI agents can fine-tune their autonomy, aligning with human team dynamics and operational needs more seamlessly [60]. LLMs have advanced conversational capabilities, enabling more natural, context-aware interactions and supporting more complex, multi-party engagements [15, 128]. Their application in mobile UI and UX evaluation also showcases their role in personalizing interactions and enhancing team collaboration [108]. The practice of chaining LLM prompts underscores their potential to customize Human-AI collaboration, ensuring greater control and transparency in the process [172].
3.1.2 Conversational Interaction
@@
In Human-AI Teams (HATs), effective communication is as crucial as in human teams, with the conversational capabilities of AI playing a pivotal role in maintaining trust, facilitating knowledge sharing, and ensuring smooth coordination. These capabilities are essential for both the cognitive processing of tasks and the emotional dynamics within the team, impacting decision-making and situational awareness [84, 135]. Quick and appropriate responses from AI can enhance human members’ confidence in their AI partners, leading to better teamwork [178].
@@
The proactive sharing of information by AI, crucial for building trust and promoting collaboration, has seen significant strides with the development of Teachable Conversational Agents. These agents improve text classification tasks by learning from dialogue-based teaching and feedback, adjusting their focus based on interaction, which can range from acquiring new knowledge to refining existing information based on the conversation’s context [25]. The perception of these agents’ likability and human-likeness is key to their effectiveness in communication [31]. Voice-based interactions, supported by Natural Language Processing (NLP), offer substantial benefits for both verbal and written exchanges in HATs [160].
@@
Research efforts have also been made to create a "Shared Vocabulary" (Taxonomy Model) for HATs, aiming for a unified set of terms and concepts accessible to both humans and AI. Furthermore, the "Communication and Explanation" framework (Communication Model) suggests equipping AI with the skills and training necessary for clear communication and explanations of their actions and reasoning, improving team understanding and efficiency [154]. This approach not only fosters mutual comprehension but significantly boosts team performance.
3.1.3 UI/UX

### v2

@@ 2.1.1. Human-AI Collaboration with Active Learning
Unlike traditional machine learning paradigms, active learning involves an iterative process where a model selectively identifies data requiring labeling, optimizing system performance with minimal training data (Freeman et al., 2014). Particularly, this approach is crucial for fine-tuning large pre-trained models, especially transformer-based ones, which require substantial labeled data for specific user scenarios. Active learning efficiently harnesses human expertise to pinpoint areas of uncertainty, enabling more targeted training with less annotation effort (Settles, 2009). The active learning workflow begins with an unlabeled dataset and a pre-trained model. The model predicts labels for each sample, outputting confidence levels (Fig. 3). When predictions fall below a quality threshold, human annotators step in for manual annotation. This iterative process of re-training the model with new labeled data continues until satisfactory confidence levels are achieved. Human AI teaming plays a vital role in efficient data labeling and continuous model improvement.
Figure 3. A visual representation of an active learning workflow. If the model predictions do not pass the required confidence value thresholds required for a task, the data is sent to humans for manual annotations, further improving the model by retraining it with the now-labeled data.
@@
Recent research in this area has introduced novel methods to leverage human expertise effectively. For instance, Pries et al. (Pries et al., 2023) propose a method for efficient data labeling using precise distance measurements to filter data samples, streamlining the presentation of data to experts. Lu et al. (Lu et al., 2022) discuss human-guided interventions for continuous model improvement, focusing on dissociating biases. Bao et al. (Bao et al., 2018) explore the transformation of human-annotated rationales into continuous attention mechanisms, enhancing the learning of domain-invariant representations. Additionally, Lertvittayakumjorn et al. (Lertvittayakumjorn et al., 2020) present a generalizable approach for debugging deep text classification models with human input, applicable to larger models. Finally, Yao et al. (Yao et al., 2023) propose an explainable-generation active learning framework for simpler tasks, potentially improving the sampling process in data augmentation.
2.1.2. Reinforcement Learning from Human Feedback
@@
Reinforcement Learning from Human Feedback (RLHF) is a paradigm that integrates human insights into the training process (Christiano et al., 2017; Ibarz et al., 2018). This approach enhances the collaboration between humans and AI systems, especially during the data control phase — a critical stage for the success of large pre-trained models. In RLHF, the human role extends beyond mere data provision to actively shaping the AI’s learning trajectory. This human-AI symbiosis is particularly effective in mitigating biases in training data, leading to fairer AI outcomes. Additionally, the human touch in RLHF makes AI decisions more transparent and comprehensible, thereby fostering greater trust among users. The adaptability of RLHF models to user preferences highlights the significance of human understanding in complex, contextual situations, enabling AI systems to be more aligned with ethical standards and personalization needs.
@@
However, the implementation of RLHF is not without its challenges. Ensuring the consistency and accuracy of human feedback, effectively scaling this approach for large datasets, and harmonizing human insights with algorithmic processes are areas that require careful consideration (Lee et al., 2023). Overcoming these challenges is essential to maximize the benefits of RLHF, paving the way for AI models that are not only technically proficient but also deeply attuned to human values and needs.
@@
Despite various theories and empirical models behind RL, as mentioned in (Kaelbling et al., 1996), RL frameworks can be abstracted as a continuous interaction between agents and their surrounding environment, in which they take actions individually or cooperatively given their states and state transitions to maximize long-term rewards. Traditionally, given a designated reward model, if the accumulated long-term reward is exactly or approximately converged under limited horizons, the agent’s behaviors and the corresponding dynamics (Sutton et al., 1999), like policy $\pi(a|s)$, Q function $Q(s,a)$, etc. are more likely to converge. Only after OpenAI ’s introduction of instructGPT (Ouyang et al., 2022) to illustrate how a reward model trained by high-quality data from labelers can optimize the LM’s ranking of generative content did research communities realize RL’s critical role in improving LM performance. Since labelers are asked to label this data in multiple turns to yield the most suitable outputs, this interaction pattern under RL formalization is considered RLHF. Furthermore, this working pattern can be widely applied in similar scenarios within feedback-acquiring processes. In the HAI teaming model, once human instructions are quantitatively converted to a reward model by ranking or scoring, the interaction between human and AI system agents can be formalized and optimized by RLHF.
@@
A notable early work in RLHF is by Ziegler et al. (Ziegler et al., 2019), where language models were trained to improve text generation and summarization based on human preferences. However, this approach, limited by its reliance on a single-run process and restricted data, faced challenges in generalization. In contrast, OpenAI’s instructGPT (Ouyang et al., 2022) introduced a three-phase training paradigm: initial training with supervised policy via demonstration data, training a reward model with labeler feedback, and optimizing the ranking policy using RL against the reward model. This methodology provides a robust baseline and ensures that GPT models evolve with continuous human feedback, leading to the creation of more sophisticated and responsive models, such as the web browser-assisted question-answering agent by Nakano et al. (Nakano et al., 2021).
Despite these advancements, RLHF remains a complex and often unstable process. To address this, Stanford’s recent research has introduced the Direct Preference Optimization (DPO) training model (Rafailov et al., 2023a). DPO simplifies policy training by directly mapping reward functions to optimal policies, circumventing the need for intricate reward maximization constraints. This approach not only surpasses traditional RLHF in controlling sentiment in generations but also enhances response quality in tasks like summarization and single-turn dialogue. Its simplicity in implementation and training offers a promising direction for further simplifying RLHF processes.
@@
The integration of Reinforcement Learning from Human Feedback (RLHF) into Human-AI (HAI) processes presents unique challenges due to the intricate steps and components involved, as evidenced by the work of Zhang et al. (Zhang et al., 2021), and Johnson and Vera (Johnson and Vera, 2019). Recognizing these complexities, recent research has shifted focus towards enhancing collaboration within the closely-coupled phases of HAI, with the aspiration of exploring further extensions in the future. A notable gap in current RLHF models is the communication barrier between AI and humans, highlighting the need for more advanced frameworks to bridge this divide.
A significant contribution to human-agent coordination is the instructRL model by Hu and Sadigh (Hu and Sadigh, 2023), which employs natural language instructions to align RL agents with human-preferred policies. This model leverages the embedded knowledge of pre-trained models, enabling AI agents to collaborate more effectively with humans across various tasks, including gaming, robotics, and control decision-making. InstructRL’s adaptability to different domains through knowledge transfer and meta-learning positions it as a versatile model for diverse HAI scenarios.
Research by Mirchandani et al. (Mirchandani et al., 2023) suggests that Large Language Models (LLMs), even without additional training, can function as proficient general sequence modelers due to their in-context learning abilities. This research explores the use of LLMs in robotics, particularly for extrapolating state sequences and prompting reward-conditioned trajectories, despite challenges like latency and compute costs. Chain-of-thought prompting further enhances LLM performance in complex reasoning tasks, as shown by Wei et al. (Wei et al., 2022). The Distillation and Retrieval of Online Corrections (DROC) system by Zha et al. (Zha et al., 2023) exemplifies the capability of LLMs to assimilate language feedback and distill knowledge from corrections, demonstrating effectiveness in novel settings.
However, McNeese et al.’s research (McNeese et al., 2021) provides a counterpoint to the notion of RLHF as a universal solution. Despite the success of NeoCITIES (McNeese et al., 2021) RL agents in simulated environments, the disparity between these environments and real-world complexities remains a significant hurdle. Furthermore, RL agents lack proficiency in non-verbal communication and still struggle with detailed verbal interactions, indicating a need for more sophisticated models that can fully comprehend both verbal and non-verbal human communication. These observations underscore the necessity for developing new HAI frameworks that go beyond existing RLHF models, calling for enhanced human integration to address the multifaceted nature of real-world environments.
2.2. Human Evaluation in AI
@@
Figure 4. Evaluation training strategies based on Clark et al. (Clark et al., 2021). They employ training methods by giving example and comparison tests, along with their correct answers. This is followed by an explanation of why the answer is correct. This way, the evaluators gradually learn about how to distinguish between human and machine-generated texts.
2.2.1. Evaluation of Trust and Transparency
@@
Establishing trust and transparency is crucial in Human-Eval-AI systems. Trust, inherently linked with transparency and understandability, aligns closely with how users perceive an AI system’s abilities and limitations. As highlighted by Bansal et al. (Bansal et al., 2019), it is essential for users to recognize the AI system’s potential for errors. This understanding enables users to develop a realistic mental model of the AI, crucial for anticipating possible failures or underperformance.
@@
Zhou and Chen’s work (Zhou and Chen, 2019) further discusses this by recommending the integration of uncertainty measures and machine learning model performance into AI-informed human decision-making processes. They advocate for a trial-to-trial approach in decision-making, where users adjust their judgment boundaries based on their experience of uncertainty and AI interactions. Their proposed Uncertainty-Performance Interface (UPI) allows users to concurrently evaluate the quality and performance of machine learning models, fostering deeper trust in AI systems. Similarly, Av et al. (AV et al., 2022) emphasize the importance of incorporating intrinsic uncertainty into the collaborative process. They suggest a mechanism for human experts to offer feedback, either as corrections to current AI recommendations or insights on preferable aspects within the search space. This feedback is integrated into a Bayesian optimization model, aiming to align the model’s inferences with the expert’s domain-specific knowledge.
2.2.2. Approaches to Improve Human Evaluation
@@
In Human-Eval-AI, human input remains essential, especially in instructing AI agents and refining AI decision-making processes. The study by Chhibber et al. (Chhibber et al., 2022) showcases the benefits of integrating human teachings into AI, not only enhancing AI performance but also building trust and influencing job delegation practices. In a similar vein, Hu and Sadigh’s instructRL (Hu and Sadigh, 2023) introduces a novel form of human-AI interaction through high-level natural language instructions, thereby improving AI interpretability and alignment with human intentions. Aaquib et al.’s research (Tabrez et al., 2020) on mental modeling in human-robot teaming reveals that mutual understanding between humans and AI significantly boosts team performance, advocating for AI systems capable of adapting to human mental models. Complementing this, Correia et al. (Correia et al., 2023) emphasize context in AI system design, proposing personalized user interactions for more relevant human-machine team actions and information.
@@
Further exploring human evaluation, Clark et al. (Clark et al., 2021) examine how human evaluators can differentiate between texts produced by humans and models like GPT-2 and GPT-3. They suggest methods to enhance evaluation accuracy, including detailed instructions, annotated examples, and text comparisons (Fig. 4). Ouyang et al. (Ouyang et al., 2022) present a different approach, fine-tuning GPT-3 into ‘InstructGPT’ using human feedback for improved user instruction alignment. Additionally, Lu et al.’s Error Analysis Prompting (EAPrompt) (Lu et al., 2023) focuses on evaluating machine translations by LLMs, leveraging Chain-of-Thought reasoning to refine LLM evaluative capabilities. Chen et al’s work (Chen et al., 2024) shows domain specific evaluation of LLM in problem-solving. Chiang and Lee (Chiang and Lee, 2023) investigate LLMs for human-like text quality assessments in specific NLP tasks, highlighting similarities with human evaluations. Finally, Uchendu et al. (Uchendu et al., 2023) highlight the effectiveness of collaborative efforts in enhancing detection capabilities, offering a unique perspective on human collaboration in evaluations.
3. Effective Human-AI Joint Systems
@@
This section concentrates on achieving efficiency in the joint human-AI (HAI) systems. The focus is on merging the strengths of different HAI systems to ensure optimal performance while maintaining compatibility between the two (Fig. 5).
3.1. Improving User Interfaces for Effective Teaming
@@
Figure 5. The chart shows the topics discussed to understand effective human-AI joint systems in this survey. The topics broadly discuss better interfacing, collaboration, and compatibility of human-AI systems.
3.1.1. Adaptation and Personalization
@@
The integration of AI agents into human teams has historically been limited by their narrow adaptability and personalization capabilities. Traditional AI systems struggled to understand human subtleties and adapt to team dynamics, being confined to specific tasks without the ability to learn or adjust, thereby reducing their effectiveness in collaborative settings (Zhang et al., 2021).
@@
The introduction of large pre-trained models has significantly revolutionized AI’s adaptability and personalization in teamwork. These models, trained on extensive datasets and utilizing sophisticated algorithms, have a markedly improved capacity to comprehend and adapt to team changes. Research demonstrates that LLM-equipped AI agents can fine-tune their autonomy, aligning with human team dynamics and operational needs more seamlessly (Hauptman et al., 2023). LLMs have advanced conversational capabilities, enabling more natural, context-aware interactions and supporting more complex, multi-party engagements (Bryan Wang, 2023; Qingxiao et al., 2023). Their application in mobile UI and UX evaluation also showcases their role in personalizing interactions and enhancing team collaboration (Mingming et al., 2023). The practice of chaining LLM prompts underscores their potential to customize Human-AI collaboration, ensuring greater control and transparency in the process (Wu et al., 2022).
3.1.2. Conversational Interaction
@@
In Human-AI Teams (HATs), effective communication is as crucial as in human teams, with the conversational capabilities of AI playing a pivotal role in maintaining trust, facilitating knowledge sharing, and ensuring smooth coordination. These capabilities are essential for both the cognitive processing of tasks and the emotional dynamics within the team, impacting decision-making and situational awareness (Kozlowski and Ilgen, 2006; Salas et al., 2017). Quick and appropriate responses from AI can enhance human members’ confidence in their AI partners, leading to better teamwork (Zhang et al., 2023).
@@
The proactive sharing of information by AI, crucial for building trust and promoting collaboration, has seen significant strides with the development of Teachable Conversational Agents. These agents improve text classification tasks by learning from dialogue-based teaching and feedback, adjusting their focus based on interaction, which can range from acquiring new knowledge to refining existing information based on the conversation’s context (Chhibber et al., 2022). The perception of these agents’ likability and human-likeness is key to their effectiveness in communication (Clark et al., 2016). Voice-based interactions, supported by Natural Language Processing (NLP), offer substantial benefits for both verbal and written exchanges in HATs (Wang et al., 2019).
@@
Research efforts have also been made to create a "Shared Vocabulary" (Taxonomy Model) for HATs, aiming for a unified set of terms and concepts accessible to both humans and AI. Furthermore, the "Communication and Explanation" framework (Communication Model) suggests equipping AI with the skills and training necessary for clear communication and explanations of their actions and reasoning, improving team understanding and efficiency (van den Bosch et al., 2019). This approach not only fosters mutual comprehension but significantly boosts team performance.
3.1.3. UI/UX